OpenStudy (anonymous):

i dont get this one , lim x--> 5 sin(x-5)/x^(2)-7x+10

ganeshie8 (ganeshie8):

hint : factor denominator

ganeshie8 (ganeshie8):

then use the sinx/x limit

OpenStudy (anonymous):

ok yea i factored the deno but how do we get sinx/x?

ganeshie8 (ganeshie8):

good question :) plugin "t = x-5" as x->5, t ->0

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x \to 5} ~\frac{\sin(x-5)}{x^2-7x+10} \) \(\large \lim \limits_{x \to 5} ~\frac{\sin(x-5)}{(x-5)(x-2)} \) let \(t = x-5\) as \(x \to 5, t \to 0\) the limit becomes : \(\large \lim \limits_{t \to 0} ~\frac{\sin(t)}{(t)(t+3)} \)

ganeshie8 (ganeshie8):

see if that looks okay so far :)

OpenStudy (anonymous):

so is the solution 1/(t+3) , right ?

ganeshie8 (ganeshie8):

nope,

ganeshie8 (ganeshie8):

\(\large \lim \limits_{t \to 0} ~\frac{\sin(t)}{(t)(t+3)} \) \(\large \lim \limits_{t \to 0} ~\left(\frac{\sin(t)}{(t)} \times \frac{1}{(t+3)}\right) \) \(\large \left(\lim \limits_{t \to 0} ~\frac{\sin(t)}{(t)}\right) \times \left( \lim \limits_{t \to 0} ~\frac{1}{(t+3)} \right)\)

ganeshie8 (ganeshie8):

take the limits

OpenStudy (anonymous):

so 1/3 right ??

ganeshie8 (ganeshie8):

\(\large \left(\lim \limits_{t \to 0} ~\frac{\sin(t)}{(t)}\right) \times \left( \lim \limits_{t \to 0} ~\frac{1}{(t+3)} \right)\) \(\large \left( 1 \right) \times \left(\frac{1}{(0+3)} \right)\)

ganeshie8 (ganeshie8):

1/3 is \(\large \color{red}{\checkmark}\)